# Optimizing Product Deduplication in E-Commerce with Multimodal Embeddings

### v1

@@ 5 Experiments & Results
In addition to vector size analysis, in Table IV the inference speed comparison with well established multimodal feature extractors are shown, all architectures here have been used in evaluation mode with the same GPU and batch size of 10, using a single worker dataloader. As can be seen, the internal model is considerably faster when inferred, which is a crucial component in a deployed environment. In addition to the infer speed, default vector output sizes for both CLIP and BLIP is far larger, making the usage of these models without modification impossible as per previous vector size feasibility results.
Building upon the vector representations, we developed a decider model designed to discern whether two products constitute a match. The model architecture, implemented using PyTorch Lightning, integrates both textual and visual features to make informed classification decisions. Specifically, the “decider” model concatenates the text and image vectors of each product pair, resulting in comprehensive feature vectors that encapsulate information from both modalities.
The architecture comprises a series of convolutional and fully connected layers. A 1D convolutional layer processes the stacked feature pairs, followed by normalization and activation to stabilize training. Subsequent linear layers further reduce the dimensionality, incorporating dropout layers to mitigate overfitting. The final output layer employs a Softmax activation function to produce probability scores for the binary classification task—determining whether a pair of products matches or not.
TABLE IV: Inference Times per 1000 Product Pairs
@@
For training the decider model, similar to image and text model we also employed the AdamW optimizer, since it is known to be performant, complemented by a learning rate scheduler that adjusts the learning rate based on validation loss improvements. The model was optimized using cross-entropy loss, aimed at minimizing classification errors. We trained the model over multiple epochs, ensuring convergence and stability in performance. The training and test datasets, consisting of 350K and 8K product pairs respectively, were obtained through an internal labeling process conducted by catalog experts. Due to the absence of a publicly available, independent test set for Turkish e-commerce product matching, our system’s performance was evaluated against an existing, black-box third-party111a commercially available API-based deduplication service solution. While this methodology introduces a known bias—as the labeled data was created from pairs generated by both our and the third-party algorithms—it served as a necessary starting point to identify potential areas for improvement and to demonstrate that our results surpassed the current solution from a business standpoint. Using the acquired test set, we compared the performance of the trained system with a third-party provider by inspecting the results of each system individually. Our system achieved 0.90 macro average F1 score and the third-party system achieved 0.83. General results of both systems can be seen in Table  V. With a better Recall performance on the match class, our system is better at detecting real duplicate products without losing precision over the third-party system.
@@
Throughout our empirical analysis, we observed specific areas where our model exhibited tendencies that warrant further investigation. Notably, the model showed a propensity to classify visually similar but distinct product variants as matches. For instance, products differing only in subtle packaging variations or minor design elements, while fundamentally representing different SKUs, were sometimes grouped together. This suggests that while our multimodal embeddings excel at capturing broad semantic and visual similarities, they occasionally struggle with fine-grained distinctions crucial for differentiating product variants that are visually very close but operationally separate. Addressing these nuances will be a key focus for future refinements.
TABLE V: Performance Comparison Results
Model Class Precision Recall F1
Internal Match 0.91 0.85 0.88
3rd Party Match 0.90 0.73 0.81
Internal Not Match 0.88 0.96 0.92
3rd Party Not Match 0.76 0.94 0.84
6 Conclusion
@@
Integrating multimodal vector representations, combining text and image embeddings, further improved duplicate detection accuracy. Our decider model, leveraging these embeddings, achieved a macro-average F1 score of 0.90, highlighting its robustness. Additionally, the use of Milvus ensures efficient similarity searches in expanding datasets, enhancing operational efficiency and customer experience. Deployed architecture is serving our internal users with 116M222as of August 2025 product vectors daily.

### v2

@@ V-C1 System Efficiency and Scalability
Beyond memory, the system is optimized for low-latency inference. The results in Table V show that our model architecture is more thant 6 times faster than CLIP and 30 times faster than BLIP when processing 1,000 product pairs. Together, these quantitative gains in memory and speed confirm that our system is not only accurate but also robust and efficient enough for a high-throughput production environment.
TABLE V: Inference Times per 1000 Product Pairs
@@
V-C2 Architectural Analysis of the Text Model
Our design process for the text model was data-driven. A preliminary analysis of 13 million product titles revealed that the most critical information is typically front-loaded, with the first 32 tokens capturing the essential details for over 95% of listings. This made a fixed 32-token input a feasible approach.
Our initial text model used a standard BERTurk backbone, where the final [CLS] token was compressed to 128 dimensions. This base model, trained on a Masked Language Modeling (MLM) task, achieved a validation loss of 1.51. However, we hypothesized that relying solely on the final, highly contextualized layer might discard valuable lower-level information from earlier layers. To capture a richer representation, we enhanced the architecture. We extracted the outputs from intermediate transformer layers—specifically layers 1, 3, 5, 7, 9, and 11—and stacked them with the output from the final layer. This multi-layer representation was then passed through a 1D convolutional chain followed by normalization and activation. This allowed the model to selectively learn the most salient features across different levels of contextualization. This enhanced model achieved a significantly lower MLM validation loss of 1.43, confirming its superior ability to capture relevant token-level information. The final [CLS] token from this improved architecture was then used as the 128-dimensional output for all downstream tasks.
V-C3 Architectural Analysis of the Image Model
The development of the image model began with a feasibility study to determine the viability of compact 128-dimensional embeddings. We trained a binary decision model using original high-dimensional representations from BERTurk and EfficientNetV2, which achieved an F1-score of 0.86. We then compressed these embeddings using PCA to various dimensions. The performance drop was minimal: 512 dimensions yielded an F1 of 0.856, 256 dimensions an F1 of 0.85, and 128 dimensions an F1 of 0.84. This minor 2% drop in F1-score for a 128-dimensional vector was a clear and acceptable trade-off for the immense gains in system efficiency.
Our first implementation was a simple AutoEncoder with an EfficientNetV2 encoder and a convolutional decoder, using Multi-scale Structural Similarity (MS-SSIM) as the loss function to avoid blurriness issues common with MSE. However, this model lacked contextual understanding; while images with similar colors and shapes had close vector representations, different variants of the same object were often mapped to distant vectors.
This led us to adopt a Masked Autoencoder (MAE) approach with a DeiT backbone to force contextual learning. To adapt MAE for e-commerce, where central details like brand logos are critical, we replaced random masking with a structured patch selection logic. For each image, we divide it into 9 patches and create a 4-patch input sequence consisting of: (1) the center patch, to preserve key details, (2) two random patches, to retain contextual understanding, and (3) a resized version of the entire image, to capture the overall object shape. The model is trained to reconstruct the full image from the resulting compressed 128-dimensional vector produced by the encoder.
V-C4 Robustness to Image Scaling
During early matching trials, we encountered a persistent issue where variations in object scale and image padding (as shown in Fig. 4) caused the vectors for the same product to be dissimilar. An initial plan to use a separate preprocessing module to remove these ”boring pixels” proved too computationally expensive for our production environment. Instead, we integrated this logic as a data augmentation step directly into the dataloader. A random number determines if the augmentation is applied. If so, the algorithm converts the image to black and white, finds the coordinates of the first non-zero pixels to form a bounding box around the object, and returns this cropped image as the model input. This process forced the model to learn scale-invariant representations, significantly improving its robustness to the diverse image styles submitted by different sellers.
V-D Error Analysis
@@
A qualitative analysis of the model’s failures revealed a primary weakness: differentiating between products with fine-grained visual distinctions. Notably, the model showed a propensity to classify visually similar but distinct product variants as matches. For instance, products differing only in subtle packaging variations or minor design elements, while fundamentally representing different SKUs, were sometimes grouped together. This suggests that while our multimodal embeddings excel at capturing broad semantic and visual similarities, they occasionally struggle with fine-grained distinctions crucial for differentiating product variants that are visually very close but operationally separate. Addressing these nuances will be a key focus for future refinements.
VI Conclusion
@@
Integrating multimodal vector representations, combining text and image embeddings, further improved duplicate detection accuracy. Our decider model, leveraging these embeddings, achieved a macro-average F1 score of 0.90, highlighting its robustness. Additionally, the use of Milvus ensures efficient similarity searches in expanding datasets, enhancing operational efficiency and customer experience. Deployed architecture is serving our internal users with 116M111as of August 2025 product vectors daily.